what is differentiation of d/dx of sin30
i have doubt that 
it is either
d/dx of cos30=-sin 30
or d/dx of cos30=0

Grade:12

1 Answers

ELITE GAMER
6 years ago
D/DX of cos30 is surely -sin30 .After which -sin30= -*1/2=-1/2D/DX cosx = -sinxD/DX since = cosx. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .. . . . . . . . .. . . . . . . . . . . . . . . . . .. . . . . . Hope u are cleared about this